OpenStudy (anonymous):

How does the length of the hypotenuse in a right triangle relate to the lengths of the legs?

OpenStudy (anonymous):

i got u

OpenStudy (anonymous):

h=hypotenuse a,b=legs Pythagorean theorem (memorize it): h2=a2+b2 solve for h: h=a2+b2−−−−−−√

OpenStudy (anonymous):

can u draw it out?

OpenStudy (anonymous):

Right Triangle Formulas Pythagorean Theorem: h^2=a^2+b^2 h=hypotenuse; a & b are legs 45° - 45° - 90° Triangle hypotenuse = side opposite 90°= leg*√2 30° - 60° - 90° Triangle leg opposite 30° = (1/2)hypotenuse leg opposite 60° = (1/2)hypotenuse*√3

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4ecc1075e4b04e045aeb4a74 go on here it is

OpenStudy (anonymous):

can i have more words and less math symbols?